OpenStudy (anonymous):

Proof by mathematical induction that (n + 1)^n < n^(n+1) for any integer n >= 3. So i've checked the base case, 64<81 but i'm not finding a way to show that (k + 1)^k < k^(k+1) implies (k + 2)^(k+1) < (k+1)^(k+2) Thanks.

OpenStudy (anonymous):

you trying show it by graphing?

OpenStudy (anonymous):

Im supposed to do it by induction, not graphing.

myininaya (myininaya):

So you know to assume \[(k+1)^k < k^{k+1}\] we want to show \[(k+2)^{k+1}<(k+1)^{k+2}\] So I guess we can try to multiply our assumption on both sides by first k+1 and see what we have then

myininaya (myininaya):

You might also find it useful after that to realize k<k+1

myininaya (myininaya):

@Juarismi are you still there?

OpenStudy (anonymous):

So\[(k+1)^{k+1}<k ^{k+2}+k ^{k+1}\] given k<k+1 then (k+1)^(k+1)>k^(k+1) How that help?

myininaya (myininaya):

Yeah I was missing the two... This problem might be a little harder than I though. You might have to rewrite the problem: Here I will help help you rewrite it: So we are supposing \[ (k+1)^k<k^{k+1} \] Now we are going to rewrite this true statement so whatever we rewrite it as we still get to assume it holds (because it is just a rewrite after all ) And I'm rewriting it because the multiplying of (k+1) on both sides wasn't very helpful. So I notice there is powers on both sides. I'm going to try to rewrite it so the power k is only on one side so we have \[(k+1)^k < k^k k => (\frac{k+1}{k})^k <k => (1+\frac{1}{k})^k < k \] Now you want to show \[\text{ if } (1+\frac{1}{k})^k <k \text{ then } (1+\frac{1}{k+1})^{k+1}<k+1 \] This should help

myininaya (myininaya):

Try to use our assumption notice in our assumption we have 1/k and then thing we want to show as 1/(k+1) but you know 1/k>1/(k+1) because 1/(k+1) gets closer to zero a lot quicker than 1/k as k goes to infinity